Find the area under the normal curve for each z-score. Shade accordingly.




1. P( < 2.20)



2. P( -2.5 < z < 2.01)




1
Expert's answer
2022-04-07T14:03:37-0400

a. Area is 0.98610 from table of z-score



b. Area is Z(2.01)-Z(-2.5)=0.97778-0.00621=0.97157